OpenStudy (anonymous):

What is the solution of the proportion? 3/4 = m/32 a. 24 b. 8 c. 3/8 d. 1/24

OpenStudy (anonymous):

24!

OpenStudy (texaschic101):

3/4 = m/32 cross multiply 4m = (32)(3) 4m = 96 m = 96/4 m = 24

OpenStudy (anonymous):

Thank you guys!

OpenStudy (texaschic101):

happy to help :)